CrazyHotGhost:

\begin{aligned} & 12x-5y = -20 \\\\ & y=x+4 \end{aligned} ​ 12x−5y=−20 y=x+4 ​ Plz help me.

rEsTiNpEaCe:

I'm assuming you want to find the point of intersection?

CrazyHotGhost:

Yes.

rEsTiNpEaCe:

So there are a couple ways of going about to find the POI (point of intersection) 1) Graphing 2) Substitution 3) Elimination Do you know how to do any of these processes?

CrazyHotGhost:

I only use the first one it comes the most naturally to me.

rEsTiNpEaCe:

Yes, you could do it by graphing on a paper and pen but it's not always the most accurate. I'd suggest you use method 2 or 3. You should learn these methods

rEsTiNpEaCe:

Why don't we do this question with substitution?

CrazyHotGhost:

Aww.... Fine

rEsTiNpEaCe:

So, as the name suggests, substitution is just substituting. We have: (1) 12x-5y = -20 (2) y = x+4 We literally just substitute eqn 2 into eqn 1. In eqn 2 it's telling us that y is equal to x+4. So for every time we see "y" in eqn 1 we'll substitute it with x+4. Give that a shot.

CrazyHotGhost:

Ok...?

rEsTiNpEaCe:

So do that and let me know what you get

CrazyHotGhost:

I got 0' 4

rEsTiNpEaCe:

Yessir

CrazyHotGhost:

Yay! Thank you.

rEsTiNpEaCe:

No problem!